# Efficient multi-prompt evaluation of LLMs

### v2

@@ 5 Assessing multi-prompt evaluation strategies
The best-prompt identification task (Shi et al., 2024) is to find the best prompt from a set of fixed templates, i.e., the one that gives the best performance for a task at hand. Shi et al. (2024) propose framing this problem as a bandit problem and using a linear model or an MLP to predict the performance of each prompt template. To apply PromptEval in this setting we use our model (3.2) and X-pIRT to estimate how good each template is coupled with sequential elimination algorithm (Azizi et al., 2021) (as in Shi et al. (2024)) to select prompt-example pairs for evaluation in each round. In Figure 4 we compare our PE to the baseline TRIPLE-GSE (Shi et al., 2024) with a logistic regression performance predictor and the same three types of covariates (PE-OneHot corresponds to PE-Rasch in previous experiments). For all covariate choices, we show that using PromptEval for best-prompt identification results in lower regret, i.e., the performance of the best template minus the performance of the chosen template. We include the full results for other benchmarks and also apply TRIPLE-GSE with an MLP in Appendix E.
@@ 6 Analysis of prompt sensitivity on MMLU
When averaged across subjects, we observe relatively small performance spreads per LLM compared to other datasets in the literature (see Figure 12 in the Appendix H). For example, we can consistently identify Llama-3-70B-Instruct as the best performing model, independent of the prompt template. On the other hand, scores within individual subjects are highly inconsistent. Figure 5 shows the distribution of prompt spreads (max-min acc.) across subjects per LLM. Most LLMs demonstrate a significant average spread of around 10% at the subject level.
@@
Within LLMs, we observe that Gemma-7B-it has a notably higher Kendall’s $W$ of 0.45 than any other model, meaning a fixed set of prompts performs best across many subjects (for full results, see Table 1 in the Appendix). Across LLMs, we do not observe high correlations within any of the subjects (see Figure 13 in Appendix H). Hence, similar to previous findings (e.g. Sclar et al., 2023), we do not identify any coherent template preferences across LLMs (for detailed results, see Appendix H).
@@ Appendix A Limitations
While our method provides a more reliable measure and a more flexible one, it assumes multiple prompts. Thus, if the question of which single prompt should be used was a challenge for old benchmarks, which set of prompt templates to use is the challenge now. While methods have been suggested for generating multiple prompts and diversifying those [Mizrahi et al., 2023], and while when many prompts are suggested the choice of each one is likely less critical, it is a limitation future work should consider.
Another limitation we note is that we do not focus on prompt engineering and do not solve this problem. While this would have been a crucial contribution for the field, we assume a set of prompts and assume we only care about evaluation and not training, which make for a useful and common setting, but it does not encompass this goal.
@@ Appendix B Adapting the correctness model for bounded $Y_{ij}$
There might be situations in LLM evaluation in which $Y_{ij}\notin\{0,1\}$ but $Y_{ij}\in[0,1]$. For example, in AlpacaEval 2.0 [Li et al., 2023], the response variable is bounded and can be translated to the interval $[0,1]$. Also, some scenarios of HELM [Liang et al., 2022] and the Open LLM Leaderboard [Beeching et al., 2023] have scores in $[0,1]$. One possible fix is changing the model for $Y_{ij}$. For example, if $Y_{ij}$ are continuous, the Beta model would be appropriate. Another possibility that offers a more immediate fix is binarizing $Y_{ij}$ as proposed by Polo et al. [2024]. That is, using a training set containing correctness data from $L$ LLMs, we could find a constant $c$ such that $\sum_{i,j,l}Y_{ijl}\approx\sum_{i,j,l}\mathds{1}[Y_{ijl}\geq c]$, where the index $l$ represents each LLM in the training set. Then, we define $\tilde{Y}_{ij}\triangleq\mathds{1}[Y_{ij}\geq c]$ and work with this newly created variable.
Appendix C Computing resources
@@
Appendix D Estimation errors by task
In Figures 6, 7, and 8, we analyze the Wasserstein 1-distance per task for each benchmark when using the method PE-EmbPT, a robust and versatile variation of PromptEval. The results show that for BBH and LMentry, the estimation errors (Wasserstein 1-distance) are more uniform across tasks compared to MMLU, where some tasks exhibit higher estimation errors. This discrepancy occurs because all tasks in BBH and LMentry have the same number of examples, whereas tasks in MMLU, particularly those with higher estimation errors, have a significantly larger number of examples when compared to the others. In those cases, a larger number of evaluations is recommended.
@@ Appendix D Estimation errors by task
Appendix E Extra results for best-prompt identification
In Figures 9, 10, and 11, we can see the full results for MMLU, BBH, and LMentry. For all benchmarks, we can see that within each triple “PE”, “GTRIPLE-SE”, “TRIPLE-MLP-GSE”, the “PE” version always has some advantage with a lower regret.
@@ Appendix E Extra results for best-prompt identification
Appendix F Theoretical results
F.1 Consistency of X-pIRT
In Theorem F.1, we claim that the X-pIRT estimator is uniformly consistent over all $i\in\mathcal{I}$.
Theorem F.1.
@@ F.1 Consistency of X-pIRT
A direct consequence of Theorem F.1 is that
@@
The proof of Theorem F.1 is embedded in the proof of Theorem 4.4.
F.2 Proof of Theorem 4.4
@@
where the third step is justified by the fact that $\sigma$ is $1/4$-Lipschitz and the last step is justified by Lemma F.2. ∎
Lemma F.4.
@@
  • If (i) holds, then there is at least $M_{2}+M_{3}+1$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s such that $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]<\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ (including $\hat{{\mathbb{E}}}[S_{i^{*}}\mid Y_{\mathcal{E}}]$). This implies that at most $M_{1}$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s will be less or equal $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]$. By Lemma F.5, we know that $j^{\prime}\neq\hat{j}$.
  • If (ii) holds, then there is at least $M_{1}+M_{2}+1$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s such that $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]>\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ (including $\hat{{\mathbb{E}}}[S_{i^{*}}\mid Y_{\mathcal{E}}]$). This implies that at most $M_{3}$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s will be greater or equal $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]$. By Lemma F.5, we know that $j^{\prime}\neq\hat{j}$.
@@
Let $\hat{i}$ and $i^{*}$ be indices in $\mathcal{I}$ such that $\hat{Q}(p)=\hat{{\mathbb{E}}}[S_{\hat{i}}\mid Y_{S}]$ and $Q(p)=S_{i^{*}}$. Notice that Lemma F.6 guarantees that $\sup_{i\in\mathcal{I}}|\hat{{\mathbb{E}}}[S_{i}\mid Y_{S}]-S_{i}|\leq\epsilon$ implies $|\hat{{\mathbb{E}}}[S_{\hat{i}}\mid Y_{S}]-S_{i^{*}}|\leq 2\epsilon$, for an arbitrary $\epsilon>0$. Consequently,
@@
holds by lemmas F.3 and F.4. Because $\epsilon>0$ is arbitrary, we have that $|\hat{Q}(p)-Q(p)|=o_{P}(1)$.
@@
Appendix G Details MMLU data
@@
Appendix H Details MMLU spread analysis
@@
Appendix I BERT fine-tuning details
I.1 Model
@@ I.3 Hyperparameters
Appendix J Heuristics for discrete features

### v3

@@ 6.2 Best-prompt identification
The best-prompt identification task (Shi et al., 2024) is to find the best prompt from a set of fixed templates, i.e., the one that gives the best performance for a task at hand. Shi et al. (2024) propose framing this problem as a bandit problem and using a linear model or an MLP to predict the performance of each prompt template. To apply PromptEval in this setting we use our model (3.2) and X-pIRT to estimate how good each template is coupled with sequential elimination algorithm (Azizi et al., 2021) (as in Shi et al. (2024)) to select prompt-example pairs for evaluation in each round. In Figure 5 we compare our PE to the baseline TRIPLE-GSE (Shi et al., 2024) with a logistic regression performance predictor and the same three types of covariates (PE-OneHot corresponds to PE-Rasch in previous experiments). For all covariate choices, we show that using PromptEval for best-prompt identification results in lower regret, i.e., the performance of the best template minus the performance of the chosen template. We include the full results for other benchmarks and also apply TRIPLE-GSE with an MLP in Appendix H.
@@ 7 Analysis of prompt sensitivity on MMLU
When averaged across subjects, we observe relatively small performance spreads per LLM compared to other datasets in the literature (see Figure 16 in the Appendix K). For example, we can consistently identify Llama-3-70B-Instruct as the best performing model, independent of the prompt template. On the other hand, scores within individual subjects are highly inconsistent. Figure 6 shows the distribution of prompt spreads (max-min acc.) across subjects per LLM. Most LLMs demonstrate a significant average spread of around 10% at the subject level.
@@
Within LLMs, we observe that Gemma-7B-it has a notably higher Kendall’s $W$ of 0.45 than any other model, meaning a fixed set of prompts performs best across many subjects (for full results, see Table 1 in the Appendix). Across LLMs, we do not observe high correlations within any of the subjects (see Figure 17 in Appendix K). Hence, similar to previous findings (e.g. Sclar et al., 2023), we do not identify any coherent template preferences across LLMs (for detailed results, see Appendix K).
@@ Appendix A Limitations
While our method offers a more reliable and flexible evaluation, it relies on using multiple prompts. As a result, if selecting a single prompt was a challenge in earlier benchmarks, determining the appropriate set of prompt templates now becomes the key challenge. Although methods have been proposed for generating and diversifying multiple prompts [Mizrahi et al., 2023], and the selection of individual prompts becomes less critical when many are used, this remains a limitation for future work to address.
Another limitation is that we do not focus on prompt engineering or attempt to solve this issue. While addressing prompt engineering would be a significant contribution to the field, our approach assumes a predefined set of prompts and focuses solely on evaluation or optimization within that pool of prompts. This is a practical and widely used setting but does not address the broader prompt engineering challenge.
@@ Appendix B Adapting the correctness model for bounded $Y_{ij}$
There might be situations in LLM evaluation in which $Y_{ij}\notin\{0,1\}$ but $Y_{ij}\in[0,1]$. For example, in AlpacaEval 2.0 [Li et al., 2023], the response variable is bounded and can be translated to the interval $[0,1]$. Also, some scenarios of HELM [Liang et al., 2022] and the Open LLM Leaderboard [Beeching et al., 2023] have scores in $[0,1]$. One possible fix is changing the model for $Y_{ij}$. For example, if $Y_{ij}$ are continuous, the Beta model would be appropriate. Another possibility that offers a more immediate fix is binarizing $Y_{ij}$ as proposed by Maia Polo et al. [2024]. That is, using a training set containing correctness data from $L$ LLMs, we could find a constant $c$ such that $\sum_{i,j,l}Y_{ijl}\approx\sum_{i,j,l}\mathds{1}[Y_{ijl}\geq c]$, where the index $l$ represents each LLM in the training set. Then, we define $\tilde{Y}_{ij}\triangleq\mathds{1}[Y_{ij}\geq c]$ and work with this newly created variable.
Appendix C Comments on the computational complexity of PromptEval
Consider the case of our experiments in which prompts are represented by embeddings of fixed size, examples are represented by one-hot encodings, and our model is given by logistic regression. Because the dimension of the embeddings does not depend on the number of prompt variations, the number of samples and variables used to fit our model does not vary with the number of prompt variations. Then, computational costs are constant with respect to the number of prompt templates. On the other hand, the number of variables (and consequently samples, to make estimation possible) should increase linearly with the number of examples, which are usually hundreds or a few thousand. Thus, this should not be a problem in most practical cases.
Appendix D Computing resources
@@
Appendix E Estimation errors by task
In Figures 7, 8, and 9, we analyze the Wasserstein 1-distance per task for each benchmark when using the method PE-EmbPT, a robust and versatile variation of PromptEval. The results show that for BBH and LMentry, the estimation errors (Wasserstein 1-distance) are more uniform across tasks compared to MMLU, where some tasks exhibit higher estimation errors. This discrepancy occurs because all tasks in BBH and LMentry have the same number of examples, whereas tasks in MMLU, particularly those with higher estimation errors, have a significantly larger number of examples when compared to the others. In those cases, a larger number of evaluations is recommended.
@@ Appendix E Estimation errors by task
Appendix F The influence of the number of prompts in PromptEval performance
We repeat the main experiment in the paper (randomly) cutting the number of prompt templates by a factor of 5. This means we use only 20 prompt variations in MMLU, for example. In summary, PromptEval still does well, beating the baseline. However, the gap between PromptEval and the baseline has shortened due to fewer variations. This fact highlights that the bigger the number of templates, the more useful PromptEval can be relative to the baseline. The results are depicted in Figure 10.
Figure 10: Performance estimation quality when we keep only 20% of the prompts variations for each task. In summary, PromptEval still does beat the baseline. However, the gap between PromptEval and the baseline has shortened due to fewer variations.
Appendix G Extra plots for the LLM-as-a-judge experiment
In Figure 11, we show the performance distribution histograms for the four considered models across prompt templates. Figure 12 shows that prompt templates consistently lead the judge to assign higher (or lower) scores across models; in this plot, we normalize the scores within rows so 0 is assigned to the lowest score and 1 is assigned to the maximum score (brighter colors denote higher scores).
Figure 11: Performance distribution histograms for the four considered models across prompt templates.
Figure 12: Prompt templates consistently lead the judge to assign higher (or lower) scores across models..
Appendix H Extra results for best-prompt identification
In Figures 13, 14, and 15, we can see the full results for MMLU, BBH, and LMentry. For all benchmarks, we can see that within each triple “PE”, “TRIPLE-GSE”, “TRIPLE-MLP-GSE”, the “PE” version always has some advantage with a lower regret.
@@ Appendix H Extra results for best-prompt identification
Appendix I Theoretical results
I.1 Consistency of X-pIRT
In Theorem I.1, we claim that the X-pIRT estimator is uniformly consistent over all $i\in\mathcal{I}$.
Theorem I.1.
@@ I.1 Consistency of X-pIRT
A direct consequence of Theorem I.1 is that
@@
The proof of Theorem I.1 is embedded in the proof of Theorem 4.4.
I.2 Proof of Theorem 4.4
@@
where the third step is justified by the fact that $\sigma$ is $1/4$-Lipschitz and the last step is justified by Lemma I.2. ∎
Lemma I.4.
@@
  • If (i) holds, then there is at least $M_{2}+M_{3}+1$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s such that $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]<\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ (including $\hat{{\mathbb{E}}}[S_{i^{*}}\mid Y_{\mathcal{E}}]$). This implies that at most $M_{1}$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s will be less or equal $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]$. By Lemma I.5, we know that $j^{\prime}\neq\hat{j}$.
  • If (ii) holds, then there is at least $M_{1}+M_{2}+1$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s such that $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]>\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$ (including $\hat{{\mathbb{E}}}[S_{i^{*}}\mid Y_{\mathcal{E}}]$). This implies that at most $M_{3}$ values of $\hat{{\mathbb{E}}}[S_{i}\mid Y_{\mathcal{E}}]$’s will be greater or equal $\hat{{\mathbb{E}}}[S_{i^{\prime}}\mid Y_{\mathcal{E}}]$. By Lemma I.5, we know that $j^{\prime}\neq\hat{j}$.
@@
Let $\hat{i}$ and $i^{*}$ be indices in $\mathcal{I}$ such that $\hat{Q}(p)=\hat{{\mathbb{E}}}[S_{\hat{i}}\mid Y_{S}]$ and $Q(p)=S_{i^{*}}$. Notice that Lemma I.6 guarantees that $\sup_{i\in\mathcal{I}}|\hat{{\mathbb{E}}}[S_{i}\mid Y_{S}]-S_{i}|\leq\epsilon$ implies $|\hat{{\mathbb{E}}}[S_{\hat{i}}\mid Y_{S}]-S_{i^{*}}|\leq 2\epsilon$, for an arbitrary $\epsilon>0$. Consequently,
@@
holds by lemmas I.3 and I.4. Because $\epsilon>0$ is arbitrary, we have that $|\hat{Q}(p)-Q(p)|=o_{P}(1)$.
@@
Appendix J Details MMLU data
@@
Appendix K Details MMLU spread analysis
@@
Appendix L BERT fine-tuning details
L.1 Model
@@
Appendix M Heuristics for discrete features